### v1

# Generative Visual Instruction Tuning

Abstract
We propose to use machine-generated instruction-following data to improve the zero-shot capabilities of a large multimodal model with additional support for generative and image editing tasks. We achieve this by curating a new multimodal instruction-following set using GPT-4V and existing datasets for image generation and editing. Using this instruction set and the existing LLaVA-Finetune instruction set for visual understanding tasks, we produce GenLLaVA, a Generative Large Language, and Visual Assistant. GenLLaVA is built through a strategy that combines three types of large pretrained models through instruction finetuning: LLama for language modeling, SigLIP for image-text matching, and StableDiffusion for text-to-image generation. Our model demonstrates visual understanding capabilities on par with LLaVA and additionally demonstrates competitive results with native multimodal models such as Unified-IO 2, paving the way for building advanced general-purpose visual assistants by effectively re-using existing multimodal models. We open-source our dataset, codebase, and model checkpoints to foster further research and application in this domain [^1] https://github.com/jeffhernandez1995/GenLlaVA.git.
1 Introduction
Large multimodal models have become increasingly popular in the research community as they are one of the key building blocks for general-purpose assistants [1, 40, 2]. The recently proposed LLaVA [25] model is among the latest wave of works that have demonstrated the effectiveness of instruction tuning for multimodal models [23, 51, 41, 14]. There has been a considerable amount of work building on top of the LLaVA model ranging from image generation [17, 36, 35], grounding [44], image editing [7] to video understanding [20]. These works share the same principles; they extend the ideas of visual instruction tuning to one or more capabilities. However, after adding a new capability (i.e., image generation), the resulting models often lose some, if not most, of their visual and language understanding capabilities. This contrasts with the original visual instruction tuning in which the models retained their language capabilities after training for visual understanding.
Figure 1: Comparison of GenLLaVA compared to recent architectures, unlike BLIP-2 [19], we use a Linear projector similar to the LlaVA architecture [25]. Generation capabilities are added using a diffusion model, but unlike GILL [17], we use a Q-former as the generation head. Finally, our model benefits from using a stronger visual encoder, namely SigLIP[49]; a stronger LLM, namely Mistral-7b [12]; and a stronger diffuser, namely SDv1.4 [33]. L stands for Linear projection, and Q stands for Q-former resampler.
In this paper, we present the generative visual instruction tunining, in which we teach a Large Multimodal Model (LMM) image understanding, image generation, and image editing tasks without significantly diminishing the performance of each individual capability. (See Fig. 1 for an overview of our method). To our knowledge, this is the first time such capability has been achieved, and our findings pave the way for building a general-purpose visual assistant. Our contributions are the following:
  • Generative multimodal instruction-following data. Inspired by previous work [25], we curate a multimodal instruction tuning set that combines image understanding, image generation, and image editing data.
  • New improved training recipe. We empirically design a single-phase training recipe to teach the LMM the necessary capabilities, unlike its predecessor, which requires multi-phase training recipes.
  • Open source. We release the following assets to the public: the generated multimodal instruction data, the codebase, the model checkpoints, and a visual chat demo.
2 Related Work
Large Multimodal Models (LMMs)
Large Multimodal Models (LMMs) refer to large language models that can understand various modalities beyond human language. Some research efforts are focused on combining image, audio, video, and other modalities with language  [50, 27], while others aim to enhance the fusion of vision knowledge and language. For example, BLIP-2 [19] created a large-scale image captioning dataset and paired a language model with a vision encoder to produce a robust multimodal model. Following this, LLaVA [25] developed a cost-effective approach to train an advanced LLM through visual instruction tuning. Although LLaVA-Next [24] improved performance for single-image tasks, it required over 2,000 tokens per image, which is about four times more than the original LLaVA. More recent models such as QwenVL[2], CogVLM[42], and Yi-VL[45] follow architectures similar to those of LLaVA. Our proposed method does not only focus on models for multimodal understanding, but also on adding generative capabilities to such models.
Generation with LMMs
In recent years, the design of methods that aim to adapt LLMs and LMMs for generative tasks has become an active research direction. For instance, GILL [17] translates the hidden representations of an LLM into embeddings that correspond to a text-to-image model by learning a neural network to perform efficient mapping using the text encoder of the diffusion model. AnyGPT [50] discretizes data from multiple modalities, extends the existing LLM vocabulary to add the extra modalities, and incorporates new randomly initialized parameters that enable additional input embeddings and prediction outputs. MGIE [7] adapts the text embedder, image input adapter, and LM head output parameters of an LMM jointly with a diffusion model for image editing from instructions. Codi-2 [38] follows a similar line and adds multiple decoders for image, audio, and video generation to condition an LLM to generate interleaved outputs from different modalities. CM3leon [46] proposes a token-based decoder-only LMM based on the CM3 architecture that is capable of both text and image generation and editing. Unified-IO 2 [27] and Chameleon [39] and GPT-4o take the early-fusion fully multimodal approach training the model from scratch to be able to expand the number of supported tasks and modalities.
3 Method
Figure 2: Editing capabilities of our model. GPT4 currently uses a version of the DALLE text-to-image model as a tool and hence is not directly able to edit images. GPT4o instead uses tools through Python generated code to accomplish the requested action. Our model, GenLLaVA, connects input features obtained from CLIP to a language model that also produces output embeddings for a text-to-image StableDiffusion model, achieving an end-to-end editing task with a multimodal model.
3.1 Background: Large Multimodal Models
Large language models (LLMs) excel in natural language generation. Large Multimodal Models enhance LLMs with the ability to interpret images and respond accordingly. Built upon a pre-trained LLM, the LMM incorporates a visual encoder (e.g., CLIP/L [31]) to derive visual features $f$, along with an adapter $\mathcal{W}$ that maps $f$ into the language domain. Following the training methodology of LLaVA [25], this process is encapsulated in the equation:
$\mathcal{C}$ =$=\{x_{1},x_{2},\ldots,x_{l}\},$ (1)
$f$ $=\text{Enc}_{\text{vis}}(\mathcal{V}),$
$x_{t}$ $=\text{LMM}(\{x_{1},\ldots,x_{t-1}\}~{}|~{}\mathcal{W}(f)),$
where $l$ represents the length of the word tokens within $\mathcal{C}$. The set $\mathcal{C}$ can represent an image caption (Features Alignment) or multimodal instruction-following data (Instruction Tuning). The LMM employs the standard autoregressive method for next-token prediction, allowing it to function as a visual assistant across diverse tasks such as visual question answering and complex reasoning. Despite gaining visual perceptive abilities through this training, its responses are currently constrained to text.
3.2 Visual Generation in Large Multimodal Models
We append $N$ visual tokens [IMG] after the instruction, with their word embeddings being trainable. The LMM learns to generate these tokens through its language modeling (LM) head. These visual tokens represent visual-related instruction comprehension within $\mathcal{E}$ and form a bridge between the language and vision modalities. We follow the same visual generation framework of GILL [17] and MGIE [7] in extracting visual features, which we summarize here for succintness.
We employ a generation head $\mathcal{T}$ to convert [IMG] into concrete visual guidance. The model $\mathcal{T}$ is a sequence-to-sequence model that translates the sequential visual tokens from the LMM into the semantically meaningful latent set $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{L}\}$ for visual guidance:
$u_{t}=\mathcal{T}(\{u_{1},\ldots,u_{t-1}\}~{}|~{}\{e_{\texttt{[IMG]}}+h_{\texttt{[IMG]}}\}),$ (2)
where $e$ denotes the word embedding and $h$ is the hidden state (from the final layer of the LMM before the LM head) of [IMG]. Specifically, the transformation applied to $e$ serves as a broad visual representation, while $h$ provides an instance-specific visual latent that reflects both the original image and the text conditioning the generation.
Table 1: Main result. Comparison of various models across advanced knowledge and general understanding. MGIE was not originally designed for these tasks, and it is purely an editing model; we take the generated caption as the answer for VQA and give a blank image as a prompt when we ask it to generate completely new images. We intend to show that models lose previous capabilities when we add a new one.
Model name Advanced Knowlege General Understanting Editing Generation
MathVista MMMU MMVet SEED-B MMB EVR CC3M COCO
GILL [17] 18.6 28.8 23.3 52.5 38.2 30.4 15.3 0.67
AnyGPT [50] 24.4 30.6 33.1 44.5 36 40.3 14.3 0.65
MGIE [7] 15.5 25.6 13 28.8 6.6 71.5 13.6 0.66
Unified-IO 2 [27] 30.7 35 36.6 61.8 71.5 50.2 13.4 0.72
GenLLaVA (Ours) 24.9 29.7 33.1 63.5 65 64.7 14.3 0.71
To guide image generation with the visual latent information $\mathcal{U}$, we employ a latent diffusion model [32], incorporating a variational autoencoder (VAE) for handling denoising diffusion in the latent space. First, we encode the desired visual output via the diffusion model encoder $o=\text{Enc}_{\text{VAE}}(\mathcal{O})$; this output may be intended for image generation or editing tasks. The diffusion process progressively introduces noise into $o$ as $z_{t}$, increasing the noise level over timesteps $t$. We then train the UNet $\epsilon_{\theta}$ to predict the added noise [10]. The diffusion process is conditioned on the visual latent information $\mathcal{U}$ through a cross-attention layer, defined as $\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{\text{dim}}})\cdot V$, where:
$Q=W^{(i)}_{Q}\cdot\varphi_{i}(z_{t}),K=W^{(i)}_{K}\cdot\{u\},V=W^{(i)}_{V}\cdot\{u\},$ (3)
with $\varphi$ representing the flattening operation, and $W^{(i)}_{Q}$, $W^{(i)}_{K}$, and $W^{(i)}_{V}$ being learnable attention matrices. We apply classifier-free guidance [11], where the score estimation $s_{\theta}$ is extrapolated to deviate from the unconditional $\varnothing$, following standard practices in diffusion models.
4 Experiment Settings
Multimodal instruction tuning is a crucial process that equips the model with a wide range of skills and capabilities across different modalities while also enabling it to adapt to novel and unique instructions. We build the multimodal instruction tuning dataset by aggregating a diverse set of supervised datasets and tasks. Each task is provided with a clear prompt, either by using existing prompts or crafting new ones using GPT4-V.
Natural Language. [1.93%]
We use the ShareGPT [34] dataset. We filter out invalid conversations, but unlike Vicuna, long conversations that surpass 2048 tokens are truncated rather than split into multiple conversations. This results in $\sim$40K conversations
Image Generation. [26.88%]
For image generation, we use the original LLaVA-Pretrain, which we now invert to be an image generation dataset instead of an image caption dataset; this dataset contains $\sim$558K data points collected from the LAION, SBU, and CC3M datasets and captioned by the BLIP-2 model.
Image Editing. [9.63%]
We subsample the IPr2Pr [3] dataset as our editing data. We use $\sim$200K CLIP-filtered data, where instructions are extracted by GPT-3, and images are synthesized by the Prompt-to-Prompt model [8].
Image Understanting.[61.56%]
We combine the original LLaVA-Finetune dataset with the LVIS-INSTRUCT4V[41], a new visual instruction tuning dataset constructed in the same way as the original LlaVA dataset but using GPT4-V as the captioner instead of BLIP-2. If they exist, we remove duplicates from the LLaVA-Finetune dataset; this results in $\sim$880K samples. We additionally add the following instruction datasets LRV-Instruction [22] ($\sim$80K), laion-gpt4v-dataset ($\sim$15K), ShareGPT4V[5] ($\sim$100K), and the following document understanding datasets DocVQA[30]($\sim$50K), SynDog-EN[16]($\sim$65K), ChartQA[29]($\sim$23K), DVQA[13] ($\sim$50K)and AI2D[15] ($\sim$15K).
4.1 Generative Visual Instruction Data
Figure 3: (Left) Results on selected Visual Question answering datasets. (Right) A qualitative example of our model.
4.2 Training details.
In this section, we evaluate our model on a broad range of tasks that require visual understanding and generation. We do not perform task-specific finetuning in any experiments. The Supplementary section details details about experimental setups, additional results, and additional studies on GenLLaVA ’s instruction capabilities.
We adopt LLaVA-v1.5-7B [23] architecture and pre-training weights, then tune it on the constructed GVIT-mix-2076K. We named this model GenLLaVA. Specifically, we replace LLaVA-v1.5-mix-665K with our GVIT-mix-2076K, add the visual generation module, and only perform a single-stage instruction tuning training as this was shown to be superior [14]. The rest of the model training protocol is kept unchanged for fair comparison. Generative Visual instruction tuning takes about 48 hours for both full-parameter tuning and LoRA tuning on 8 NVIDIA Tesla A100 GPUs, each with 80GB memory, with DeepSpeed ZeRO Stage 3.
Visual Understanting
We evaluate vision language performance and compare it against other generalist models, i.e., models that can do visual generation and understating. Results on a collection of 8 vision/language benchmarks are shown in Table 1. These benchmarks are designed to probe advanced knowledge and general understanding. MMBench [26] evaluates answer robustness through comprehensive shuffling of multiple-choice options. SEED-Bench [18] tests model performance on images and videos with multiple-choice questions. MM-Vet [47] evaluates the ability to engage in visual conversations and assess the accuracy and helpfulness of responses. Mathvista [28] consolidates mathematical reasoning benchmarks, focusing on logical and algebraic reasoning with puzzle tests. MMMU [48] covers 57 subjects across STEM, humanities, and social sciences, ranging from elementary to advanced professional levels, and tests both world knowledge and problem-solving skills.
Visual generation.
For visual generation evaluation, we measured the Fréchet Inception Distance (FID) [9] on the CC3M validation set [4] (which is a measure of image realism) and CLIP Similarity on the MS-COCO dataset [21] (which is a measure of the alignment between the text prompt and the generated image). For image editing, we measured the DINOScore on the 5.7K examples of the EVR dataset [37] following the same protocol of Fu et.al. [7].
Figure 4: Ablation on the effect of adding extra data to our instruction set.
4.3 Main Result
We compare our model, GenLLaVA, against state-of-the-art models on benchmarks for advanced knowledge, general understanding, visual generation, and editing.
As shown in Table 1, GenLLaVA consistently outperforms previous models (GILL, AnyGPT, and MGIE) in most tasks. It achieves higher scores on MathVista (24.9 vs. GILL’s 18.6, AnyGPT’s 24.4, MGIE’s 15.5, Unified-IO 2’s 30.7) and MMMU (29.7 vs. GILL’s 28.8, AnyGPT’s 30.6, MGIE’s 25.6, Unified-IO 2’s 35). For general understanding, GenLLaVA excels in SEED-B (63.5 vs. Unified-IO 2’s 61.8, GILL’s 52.5, AnyGPT’s 44.5, MGIE’s 28.8) and MMB (65 vs. Unified-IO 2’s 71.5, GILL’s 38.2, AnyGPT’s 36, MGIE’s 6.6).
In image editing on the EVR dataset, GenLLaVA scores 64.7, behind MGIE’s 71.5 but ahead of Unified-IO 2 (50.2), GILL (30.4), and AnyGPT (40.3). For visual generation, GenLLaVA has a FID score of 14.3 on CC3M, close to Unified-IO 2 (13.4) and matching AnyGPT. On the CLIP Similarity metric for MS-COCO, GenLLaVA scores 0.71, close to Unified-IO 2’s 0.72 and ahead of GILL (0.67), AnyGPT (0.65), and MGIE (0.66).
While GenLLaVA does not surpass Unified-IO 2 in all categories, it shows strong, balanced performance across various tasks, highlighting its versatility and robustness as a generalist vision-language model with less training time.
4.4 Results on selected Visual Question Answering datasets.
We evaluate various models’ performance across diverse visual question-answering datasets, including VQAv2, GQA, VisWiz, TextVQA, and ScienceQA. Our results show that Unified-IO 2 and GenLLaVA consistently perform well across most datasets. Specifically, Unified-IO 2 achieves the highest scores on the ScienceQA (86.2%) and TextVQA (67%), while GenLLaVA demonstrates strong performance on VQAv2 (79.3%) and a competitive score on GQA (62.9%). In contrast, GILL and MGIE exhibit generally lower performance across all datasets, with MGIE notably struggling on VisWiz (20.1%) and TextVQA (26.3%). AnyGPT shows moderate effectiveness, with its best performance on ScienceQA (61%). We used VLMEvalKit [6] to get the results for these datasets, which perform a generation-based evaluation using an LLM judge222 We used GPT-4 (0409) as the judge., the results can be seen in Fig. 3.
4.5 Ablations
We investigate the effect of scaling the data used to train GenLlaVA, the effect of using different image backbones, and the number of visual tokens used for image generation.
Instruction data.
We start with the original instruction tuning dataset from LlaVA-1.5, basically reproducing the original results using a one-stage training recipe. We add the textttLLaVA-Pretrain and generation head to our model; we notice that adding generation capabilities significantly affects the visual understanding capabilities in the model where all the metrics between 2.3% (MM-Vet) to 7.9% (MathVista) points in performance, to compensate for this loss in performance we modify the ration of image generation to image understanding data in our dataset from approximately $\sim$50%-50to $\sim$70%-30%, by adding more image instruction data from LVIS-INSTRUCT4V, LRV-Instruction and other chart understanding datasets. This results in a model with significant generation capabilities that maintain its image-understanding capabilities. We finally add image-generation capabilities using our selected subset of the IPr2IPr dataset. This reduces the image understanding capabilities, but we consider this a small enough change that balances the three tasks while maintaining commendable performance. The results can be seen in Fig. 4a.
Choice of Visual encoder.
The quality of the vision encoder can also have big effects on the final LMM performance. We start by using a CLIP/B, which has the lowest performance, then we compare against a stronger visual encoder and create versions of GenLLaVA, which is trained with CLIP [31] and SigLIP [49], respectively. We can see in Fig. 4b that the SigLIP encoder can achieve generally better performance CLIP encoder. This shows that SigLIP is a better vision encoder for LMM development.
Number of visual generation tokens.
We experiment with varying the number of visual generation tokens, $N$. We noticed that we need more visual generation tokens than GILL ($N=4$) and MGIE ($N=8$) to achieve the best performance. We find that $N=16$ is the best choice for our model. We hypothesize that this is because our model has to balance image generation and editing in the same head and thus needs more visual tokens to capture the complexity of the tasks. The results can be seen in Fig. 3c.
Figure 5: Qualitative conversational example of our model. The dashed line indicates that the conversation has to be restarted from the beginning due to the model losing track of it.
4.6 Comparision with SOTA.
When compared with state-of-the-art models, GenLLaVA maintains similar performance to models of the LlaVA family when evaluated using the average of the scores on the MathVista, MMMU, MMVet, SEED-B, and MMB datasets but lags behind bigger and more specialized models. However, some of these models lack the generative capabilities present in GenLLaVA.
When compared with models of similar size and setup ($\sim 7$b parameters), our model surpasses the original LlaVAv1[25] model by  5% points (37.5% vs. 43.2%), but lags behind LlaVA-1.5 by $\sim$2points (45.3% vs. 43.2%). It is surpassed by the LlaVA-Next [24] family of models by $\sim$7points; by Idefics2 by $\sim$12points (55.7% vs. 43.2%) and by the newly released MiniCPM-Llama3 [43] model by $\sim$17points (60.6% vs. 43.2%). Compared with the absolute state-of-the-art open-source models, GenLLaVA lags behind Yi-VL [45] by $\sim$6points; Emu2 [35] by $\sim$2points, LlaVA-Next [24] (34b) by $\sim$15points; and InternVL 1.5 by $\sim$20points. Compared with the absolute state-of-the-art closed models, GenLLaVA lags behind GPT-4o by $\sim$25points, GPT-4V [1] by $\sim$24points; and the Gemini family [40] of models by  16% points.
5 Conclusion
In this paper, we have introduced the Generative Large Language, and Visual Assistant (GenLLaVA), a new approach that enables Large Multimodal Models (LMMs) to simultaneously excel in image understanding, generation, and editing without significant loss in individual task performance. We balanced multimodal capabilities in a single model by curating a comprehensive multimodal instruction dataset and developing an innovative single-phase training methodology. Our work sets a new standard for building visual assistants with extra capabilities, and we hope our open-source contributions, including datasets, codebase, and model checkpoints, will serve as valuable resources for the research community, driving further advancements in the field of multimodal AI.
Acknowledgements
The authors would like to thank Google Cloud and the TPU Research Cloud program from Google for providing funding for this research effort. We are also thankful for support from the Department of Computer Science at Rice University.